OpenStudy (anonymous):

A certain highway undergoes a difference in elevation of 1600 feet in 8 miles. What is the average percent slop of the highway's grade?

myininaya (myininaya):

do you know what slope is?

myininaya (myininaya):

it is a change of something over another change of something

OpenStudy (anonymous):

I do not. That's the question. In the textbook, a similar question is asked and its is "1800 feet and 7 miles" and the percentage is 5.

myininaya (myininaya):

you are given two somethings here feet and miles use that in my definition of a slope

OpenStudy (anonymous):

I don't get how to set it up to get a percentage

myininaya (myininaya):

make sure you also write both things in the same unit

myininaya (myininaya):

set the fraction up first

myininaya (myininaya):

then we will write the fraction as a percentage by just dividing

OpenStudy (anonymous):

so 8 miles in feet is 42240

OpenStudy (anonymous):

so 1600/42240 = .0378...

myininaya (myininaya):

right which you can write .0378... has a percentage just like you would do for that other problem you put which i will do here so you can have an example \[\frac{1800ft}{7m}=\frac{1800 ft}{7 m} \cdot \frac{1 m }{ 5280 ft} =\frac{1800}{7(5280)}=\frac{1800}{36960} \approx 0.0478 =4.78 \%\]

myininaya (myininaya):

which your book rounded to 5 percent

OpenStudy (anonymous):

AHHH now I see it. It has been too long since I've done math like this. It's so easy that it's hard. Thanks so so much!

myininaya (myininaya):

lol np

myininaya (myininaya):

so you can say your answer is what?

myininaya (myininaya):

write .0378 as a percent

myininaya (myininaya):

remember I wrote .0478 as 4.78 percent basically the decimal point was just moved over twice